OpenStudy (anonymous):

Chapter 5:Indices and Logarithms

OpenStudy (anonymous):

\[Given~\log _{3}x=7~and~\log _{3}y=3,\]find the values of:\[\log _{3}27x^2y\]

OpenStudy (anonymous):

My answer is 19 but the answer in the book say it's 20

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@iambatman

Nnesha (nnesha):

first use rules to expand

Nnesha (nnesha):

multiplication .---> addition division ----- > subtraction

OpenStudy (anonymous):

\[\log _{3}3^2+\log _{3}x^2+\log _{3}y\]

OpenStudy (anonymous):

\[\log_3 27x^2y = \log_3 27 + 2\log_3 x + \log_3 y = 3+2(7) +3 = 20\]

Nnesha (nnesha):